# BotArtist: Generic approach for bot detection in Twitter via semi-automatic machine learning pipeline

### v4

@@ 5.1 Models comparison
6 Conclusions and Future Work
In this research paper, we introduce a semi-automatic machine learning pipeline (SAMLP) that addresses multiple challenges in creating machine learning models, including feature selection, hyperparameter fine-tuning, model evaluation, decision threshold optimization for binary classification, and provides model explainability through the SHAP game-theoretical approach. We apply this developed pipeline to create BotArtist, a versatile bot detection model based on user profile features. Our approach is trained and evaluated alongside current state-of-the-art solutions using nine different datasets. As demonstrated in our experiments, BotArtist surpasses existing, more complex methods in terms of generalization, achieving almost a 10% increase in the total F1 score in comparisons of multiple data sets and a 6. 5% increase in comparisons of separated data sets. Additionally, we offer insights into the final model’s decision-making process and the patterns it captures, based on the SHAP game-theoretical approach (available at the repository). One of the limitations of the presented methodology is the relatively limited feature set, which might make it susceptible to evasion by future bot accounts that specifically aim to avoid detection based on these features. Further research conducted over a longer period is required to assess the model’s effectiveness in detecting bot accounts in the evolving landscape of social networks.
@@ 6 Conclusions and Future Work
Appendix A Appendix
A.1 Model benchmarks
As mentioned in section 5 we compare BotArtist implementation with the existing 35 bot detection methods. We replicate the experiments of the model benchmarks described in (Feng et al., 2022a) where the authors provide the implementation of the 34 models. The Botometer (Yang et al., 2022) is an available model on-line that does not require additional training. Furthermore, the authors provide the original dataset split for training, validation, and testing portions, which allows the replication of the experiments with unbiased configurations. We manage to re-train each model (except the Botometer) over the training/validation set and evaluate their performance over the test set, as provided by (Feng et al., 2022a) (see Table 3).
We differentiate from the original paper in the generic case scenario, as we combine the training/validation portions of the nine datasets into a single train portion where the selected top-performance models were trained over the combination of the datasets to evaluate the generalization of the methods. To reproduce the experiments, the repository 444GitHub repository: TBD contains all the necessary scripts and information. The designed experiments were executed on a Dell server machine with two AMD EPYC 7543 CPUs (32 cores/64 threads), 512GB DDR4 RAM (16X32 GB) and nVidia A30 GPUs (24 GB VRAM), with a total execution time around 30 days.
A.2 Created dataset
#Ukraine, #Ukraina, #ukraina, #Украина(Ukraine), #Украине(Ukraine), #PrayForUkraine, #UkraineRussie, #StandWithUkraine, #StandWithUkraineNOW, #RussiaUkraineConflict, #RussiaUkraineCrisis, #RussiaInvadedUkraine, #WWIII, #worldwar3, #Война(War), #BlockPutinWallets, #UkraineRussiaWar, #Putin, #Russia, #Россия(Russia), #StopPutin, #StopRussianAggression, #StopRussia, #Ukraine_Russia, #Russian_Ukrainian, #SWIFT, #NATO, #FuckPutin, #solidarityWithUkraine, #PutinWarCriminal, #PutinHitler, #BoycottRussia, #with_russia, #FUCK_NATO, #ЯпротивВойны(I’m against war), #StopNazism #myfriendPutin #UnitedAgainstUkraine #StopWar #ВпередРоссия(Go Russia), #ЯМыРоссия(I/we Russia), #ВеликаяРоссия(Great Russia), #Путинмойпрезидент(Putin is my president), #россиявперед(Go Russia), #россиявперёд(Go Russia), #ПутинНашПрезидент(Putin is our president), #ЗаПутина(For Putin), #ПутинВведиВойска(Putin send the troops), #СЛАВАРОССИИ(Glory to Russia), #СЛАВАВДВ(Glory to Russian Airborne Forces) |
The provided 10M Twitter profiles dataset was retrieved through the Twitter API, while still available to the research community. The collection of the user profiles was initiated at the beginning of the 2022 Russo-Ukrainian War on 24 February 2022. We identify most of the popular hashtags based on the Ukrainian, Russian, and English hashtags during the initial days of invasion (see Table 5). Based on the collected hashtags we initiate the collection of the respective user profiles using them in their posts. Additionally, by using Twitter API we collect data from previous periods (23 February 2022) to capture some profiles and their activity 24 hours before the invasion. Furthermore, we collect on a daily basis the user profiles for the 16-month period until June 23, 2023.
To share the collected user profiles and not violate user privacy and rights of the removal of their information from online social media, we provide the dataset in the processed form of CSV files where only numerical features are presented without any uuid information such as username, screen_name, description, URLs, user_id and profile_image_url. To hide the private user data while providing as much useful information for further usage in the case of ML/AI applications, we compute as many features as possible to maximize the information provided by the dataset (see Table 2).
Additionally, based on the described BotArtist model we predict the collected user profiles and reveal which profile belongs to a human or bot account, based on BotArtist in combination with the model probability. Furthermore, we contact the original authors of the existing anonymized textual dataset over the same topic of Russo-Ukrainian War (Shevtsov et al., 2022a). We provide identical non-reversible anonymized uids to keep the connection between the user profile information and their tweets or user mentions over the 127.275.386 tweets. We believe that the combination of the provided labels and the large volume of textual data in the new era of Large Language Models will help the research community to develop state-of-the-art solutions for bot content detection. For this purpose the provided Zenodo dataset 555Zenodo data repository: TBD and the metadata 666GitHub repository: TBD are shared over the Attribution-NonCommercial-ShareAlike 4.0 International license (CC BY-NC-SA 4.0 DEED) 777Dataset license: https://creativecommons.org/licenses/by-nc-sa/4.0/. Shared data will not be maintained due to updated Twitter API subscriptions and will be provided only in the current public version of Zenodo repository in form of CSV file (see data sample provided in Table 6 ).
A.3 Explainability
This study aims to create a high performance model that can be used for bot detection while shedding light on the decision of the model itself. For this purpose, we employ SHapley Additive exPlanations (SHAP) values (Lundberg and Lee, 2017) which offer various benefits. One of the most significant advantages of SHAP values is that they are not tied to any particular type of machine learning model and, therefore, are model-agnostic. Furthermore, SHAP values display local accuracy and consistency and cannot be found concurrently in other methods.
In this section, we explain the BotArtist model trained over multiple datasets. Figure: 2 presents the 20 most important features and their impact on the final decision. The features are sorted in descending order of importance. Based on this order, we observe that the age of the account and the level of social activity are the most critical characteristics of our model, since the creation date of the account.
The presented figure showcases the values of the features, ranging from low (indicated by the blue color) to high (indicated by the red color). The x-axis provides information on the impact of each feature value on the model decision. Based on the results presented, it is crucial to note that the SHAP explanation technique provides insights that align with past related work. For instance, we reveal that bot accounts tend to have a significantly lower age of the account and a lower number of followers combined with a higher number of followers by age of the account. Based on this information, we can assume that bot accounts are not so popular in terms of total followers in comparison with normal users, while the growth of the profile followers is higher in comparison with normal users. This assumption agrees with the original target of the bot account creators to achieve a higher audience during the short period.

### v5

@@ Introduction
Conclusions and Future Work
In this research paper, we introduce a semi-automatic machine learning pipeline (SAMLP) that addresses multiple challenges in creating machine learning models, including feature selection, hyperparameter fine-tuning, model evaluation, decision threshold optimization for binary classification, and provides model explainability through the SHAP game-theoretical approach. We apply this developed pipeline to create BotArtist, a versatile bot detection model based on user profile features. Our approach is trained and evaluated alongside current state-of-the-art solutions using nine different datasets. As demonstrated in our experiments, BotArtist surpasses existing, more complex methods in terms of generalization, achieving almost a 10% increase in the total F1 score in comparisons of multiple data sets and a 6. 5% increase in comparisons of separated data sets. Additionally, we offer insights into the final model’s decision-making process and the patterns it captures, based on the SHAP game-theoretical approach (available at the repository).
One of the limitations of the presented methodology is the relatively limited feature set, which might make it susceptible to evasion by future bot accounts that specifically aim to avoid detection based on these features. Further research conducted over a longer period is required to assess the model’s effectiveness in detecting bot accounts in the evolving landscape of social networks.